OpenStudy (anonymous):

How would you solve How would you solve √(2x+5)=2√2 +1?

OpenStudy (danjs):

like this like this \[[\sqrt{x}]^2=x\]

OpenStudy (danjs):

\[[\sqrt{2x+5}]^2=[2\sqrt{2}+1]^2\]

OpenStudy (anonymous):

Okay... so lemme try this

OpenStudy (anonymous):

Oh there should be an "x" after the second √2

OpenStudy (danjs):

you didn't type one, idk, look at the prob

OpenStudy (anonymous):

Yeah, I meant I mistyped it haha

OpenStudy (anonymous):

oh I see, I just didn't foil it when I did it last time... thanks!

OpenStudy (anonymous):

Appreciate it :).

OpenStudy (danjs):

\[[2x+5 = (2\sqrt{2}+1)(2\sqrt{2}+1)\]

OpenStudy (danjs):

welcome

OpenStudy (anonymous):

My textbook only lists one answer. Shouldn't there be two?

OpenStudy (danjs):

what is the prob, didnt you say you typed it in wrong

OpenStudy (anonymous):

Pretty much the second thing you typed

OpenStudy (anonymous):

except the 2x+5 is square root

OpenStudy (danjs):

\[\sqrt{2x+5}=2\sqrt{2}+1\] that?

OpenStudy (anonymous):

The discriminant is positive so there should be two right? And yes, but after the root 2 on the right there should be an x (so root(2x))

OpenStudy (danjs):

\[2x+5=(2\sqrt{2}*x+1)(2\sqrt{2}*x+1)\]

OpenStudy (anonymous):

Yes, and I got the correct answer

OpenStudy (anonymous):

But only 1 listed in the textbook whereas I got two

OpenStudy (danjs):

\[2x+5=8x^2+4\sqrt{2}*x+1\]

OpenStudy (anonymous):

Yep got the right answer using that

OpenStudy (anonymous):

There should be two though right?

OpenStudy (danjs):

only the positive answer

OpenStudy (anonymous):

both are positive... hmm whatever! I guess maybe the txtbook just ommitted one of them

OpenStudy (anonymous):

If you don't mind, how would I solve x^2-6ix-9?

OpenStudy (anonymous):

Would I just.. factor?

OpenStudy (danjs):

you cant solve that, there isn't an equation

OpenStudy (anonymous):

I think they want the imaginary answers

OpenStudy (anonymous):

ooohhh quadratic formula nvm

OpenStudy (danjs):

what do the directions say, it shouldn't be Solve

OpenStudy (anonymous):

Kk thanks!